# Emergent classical gravity as stress-tensor deformed field theories

### v1

@@ A. Exact deformed action from perturbation method
$S$ $=\hat{S}[\hat{\gamma}+h,\psi]+\frac{1}{2\lambda}S_{\text{grav}}[\hat{\gamma}+h]$
$=\hat{S}[\hat{\gamma},\psi]-\frac{1}{2}\int d^{d}x\sqrt{\hat{\gamma}}h_{\mu\nu}\hat{T}^{\mu\nu}-\frac{1}{2}\sum_{k=1}^{\infty}\int d^{d}xh_{\mu\nu}\Big{(}\prod_{i=1}^{k}h_{\mu_{i}\nu_{i}}\Big{)}\frac{\partial^{k}}{\prod_{i=1}^{k}\partial\hat{\gamma}_{\mu_{i}\nu_{i}}}\Big{(}\sqrt{\hat{\gamma}}\hat{T}^{\mu\nu}\Big{)}$
@@
$\mathcal{F}^{(k)}$ $=\mathcal{F}^{(k)}(\hat{\gamma}_{\mu\nu},h_{\mu\nu},\hat{\nabla}_{\rho}h_{\mu\nu},\hat{\nabla}_{\rho}\hat{\nabla}_{\sigma}h_{\mu\nu})\sim h^{k},$
@@
The EOM for auxiliary field is
$-\lambda l^{-2}\Big{[}{\hat{T}^{\mu\nu}+\sum_{k=1}^{\infty}\tau^{k}\Big{(}\prod_{i=1}^{k}h_{\mu_{i}\nu_{i}}\Big{)}\frac{1}{\sqrt{\hat{\gamma}}}\Big{(}\frac{\partial^{k}}{\prod_{i=1}^{k}\partial\hat{\gamma}_{\mu_{i}\nu_{i}}}(\sqrt{\hat{\gamma}}\hat{T}^{\mu\nu})+k\frac{\partial^{k}}{\partial\hat{\gamma}_{\mu\nu}\prod_{i=2}^{k}\partial\hat{\gamma}_{\mu_{i}\nu_{i}}}(\sqrt{\hat{\gamma}}\hat{T}^{\mu_{1}\nu_{1}})\Big{)}}\Big{]}$
$\quad+\sum_{k=0}^{\infty}\tau^{k}\tilde{\mathcal{F}}^{(k)\mu\nu}=0,$ (14)
@@
$\tilde{\mathcal{F}}^{(k)\mu\nu}(x)=\frac{1}{\sqrt{\hat{\gamma}(x)}}\frac{\delta}{\delta h_{\mu\nu}(x)}\int d^{d}x^{\prime}\sqrt{\hat{\gamma}}\mathcal{F}^{(k+1)}\sim h^{k}.$ (15)
@@
$\tilde{\mathcal{F}}^{(0)\mu\nu}$ $=\frac{1}{2}(\hat{R}-2\Lambda)\hat{\gamma}^{\mu\nu}-\hat{R}^{\mu\nu},$
$\tilde{\mathcal{F}}^{(1)\mu\nu}$ $=\frac{1}{2}\hat{\Box}h^{\mu\nu}-\frac{1}{4}\hat{\gamma}^{\mu\nu}\hat{\Box}h^{\alpha}_{\alpha}+\frac{1}{2}\Big{[}\hat{R}^{\mu\rho\nu\sigma}+\hat{R}^{\nu\rho\mu\sigma}+\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}+\hat{R}^{\nu\rho}\hat{\gamma}^{\mu\sigma}$
$\quad-\hat{R}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-\hat{R}^{\rho\sigma}\hat{\gamma}^{\mu\nu}-\frac{1}{2}(\hat{R}-2\Lambda)(2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\Big{]}h_{\rho\sigma}.$ (16)
@@
$h^{*}_{[k]\mu\nu}=0,\text{ for }k\leq 0.$ (18)
Such a saddle point corresponds to the seed theory with a non-local stress tensor deformation, and the form of the deformation can be solved order by order using perturbation method. Below, we proceed to calculate the deformation up to the second order in $\lambda$. One can easily find that
$\tilde{\mathcal{F}}^{(2)\mu\nu}$ $=\frac{1}{8}\Big{(}4\hat{\nabla}_{\rho}h^{\mu\nu}\hat{\nabla}_{\sigma}h^{\rho\sigma}-8\hat{\nabla}_{\rho}h^{\mu\rho}\hat{\nabla}_{\sigma}h^{\nu\sigma}-2h^{\mu\nu}\hat{\nabla}_{\sigma}\hat{\nabla}_{\rho}h^{\rho\sigma}+8h^{\nu\rho}\hat{\nabla}_{\sigma}\hat{\nabla}_{\rho}h^{\mu\sigma}+4h^{\rho\sigma}\hat{\nabla}_{\sigma}\hat{\nabla}_{\rho}h^{\mu\nu}+2h^{\mu\nu}\hat{\Box}h^{\rho}_{\rho}$
@@
$\tau h^{*}_{\mu\nu}(x)$ $=-2\lambda t_{\mu\nu}(x)+4\lambda^{2}l^{-2}\int d^{d}y\sqrt{\hat{\gamma}(y)}G_{\mu\nu\eta\xi}(x,y)\Big{(}\delta^{\eta}_{\alpha}\delta^{\xi}_{\beta}-\frac{1}{D-2}\hat{\gamma}^{\eta\xi}(y)\hat{\gamma}_{\alpha\beta}(y)\Big{)}\Big{[}M^{\alpha\beta\rho\sigma}t_{\rho\sigma}$

### v2

@@ A. Exact deformed action from perturbation method
As an illustrated example, we consider the first-order non-local stress tensor deformation of the free scalar field theory when the background metric is flat,
$S^{(\lambda)}[\eta,\phi]$ $=\hat{S}[\eta,\phi]+\frac{\lambda l^{-2}}{2}\int d^{d}x\sqrt{\eta}\hat{T}_{\mu\nu}\frac{1}{-\hat{\Box}}(\hat{T}^{\mu\nu}-\frac{1}{d-2}\hat{T}^{\alpha}_{\alpha}\eta^{\mu\nu})+O(\lambda^{2})$
$=\frac{1}{2}\int d^{d}x\sqrt{\eta}\Big[\eta^{\rho\sigma}\partial_{\rho}\phi\partial_{\sigma}\phi+\lambda l^{-2}\big(\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}\eta_{\mu\nu}\eta^{\rho\sigma}\partial_{\rho}\phi\partial_{\sigma}\phi\big)\frac{1}{-\hat{\Box}}\partial^{\mu}\phi\partial^{\nu}\phi\Big]+O(\lambda^{2}).$ (13)
By introducing the metric redefinition
$g_{\mu\nu}$ $=\eta_{\mu\nu}-2\lambda l^{-2}\frac{1}{-\hat{\Box}}\partial_{\mu}\phi\partial_{\nu}\phi+O(\lambda^{2}),$
$g^{\mu\nu}$ $=\eta^{\mu\nu}+2\lambda l^{-2}\frac{1}{-\hat{\Box}}\partial^{\mu}\phi\partial^{\nu}\phi+O(\lambda^{2}),$ (14)
and plugging them into the first order deformed action (Geometric realization of stress-tensor deformed field theory), we obtain
$S^{(\lambda)}$ $=\frac{1}{2}\int d^{d}x\sqrt{g}g^{\rho\sigma}\partial_{\rho}\phi\partial_{\sigma}\phi-\frac{l^{2}}{8\lambda}\int d^{d}x\sqrt{\eta}\Big((g^{\mu\nu}-\eta^{\mu\nu})\hat{\Box}(g_{\mu\nu}-\eta_{\mu\nu})-\frac{1}{2}(\eta_{\mu\nu}g^{\mu\nu}-d)\hat{\Box}(\eta^{\rho\sigma}g_{\rho\sigma}-d)\Big)+O(\lambda^{2}),$ (15)
which can be regarded as the free scalar field theory on the metric $g$ coupled with the linearized Einstein’s gravity.
A further illustrative example is that of Einstein-Maxwell theory, whose gravitational action is given by
$S_{\text{grav}}^{(\lambda)}[g,A]=-\frac{1}{4}\int d^{4}x\sqrt{-g}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2\lambda}\int d^{4}x\sqrt{-g}R.$ (16)
This system is first analyzed using the method of the effective deformed action introduced in the main text. Taking the variation with respect to the metric, we find the Einstein’s equation
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=2\lambda\hat{T}_{\mu\nu},$ (17)
where
$\hat{T}_{\mu\nu}=F_{\mu\alpha}F_{\nu}{}^{\alpha}-\frac{1}{4}g_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta}.$ (18)
The metric solution is denoted as $g^{(\lambda)}$. Taking the trace of the Einstein’s equation, and using the fact that $g^{\mu\nu}\hat{T}_{\mu\nu}=0$, we find
$R(g^{(\lambda)})=\hat{T}(g^{(\lambda)})=0.$ (19)
Consequently, the effective deformed theory retains the same functional form as Maxwell’s theory, while the metric is replaced by the solution to the Einstein’s equation.
$S_{\text{EDA}}[g^{(\lambda)},A]=\Big(S_{\text{Maxwell}}[g,A]+S^{(\lambda)}_{\text{EH}}[g]\Big)\Big|_{g=g^{(\lambda)}}=S_{\text{Maxwell}}[g^{(\lambda)},A].$ (20)
We can further prove that when the metric satisfies the Einstein’s equation, the effective deformed action is independent of the deformation parameter $\lambda$, i.e.,
$\frac{\text{d}}{\text{d}\lambda}S_{\text{EDA}}[g^{(\lambda)},A]=\frac{\text{d}}{\text{d}\lambda}S_{\text{Maxwell}}[g^{(\lambda)},A]=0.$ (21)
There are two approaches to prove this identity. The first is to consider the derivative of the gravitational action $S^{(\lambda)}[g^{(\lambda)},A]$ with respect to the deformation parameter,
$\frac{\text{d}}{\text{d}\lambda}S_{\text{EDA}}[g^{(\lambda)},A]$ $=\frac{\text{d}}{\text{d}\lambda}\Big(S_{\text{Maxwell}}[g,A]+S^{(\lambda)}_{\text{EH}}[g]\Big)_{g=g^{(\lambda)}}$
$=\frac{\partial}{\partial\lambda}S^{(\lambda)}_{\text{EH}}[g^{(\lambda)}]+\int d^{4}x\frac{\delta(S_{\text{Maxwell}}+S^{(\lambda)}_{\text{EH}})}{\delta g^{\mu\nu}}\Big|_{g=g^{(\lambda)}}\frac{\partial g^{(\lambda)\mu\nu}}{\partial\lambda}$
$=-\frac{1}{2\lambda^{2}}\int d^{4}x\sqrt{-g^{(\lambda)}}R(g^{(\lambda)})=0.$ (22)
The second approach examines the action of Maxwell theory on the metric saddle $g^{(\lambda)}$. Since the stress tensor of Maxwell theory is traceless, this yields an integral identity
$\int d^{4}x\sqrt{-g^{(\lambda)}}R(g^{(\lambda)})=0.$ (23)
Consider a small variation of the deformation parameter, $\lambda\to\lambda+\delta\lambda$, we find
$\delta\int d^{4}x\sqrt{-g^{(\lambda)}}R(g^{(\lambda)})=\int d^{4}x\sqrt{-g^{(\lambda)}}\Big(R_{\mu\nu}(g^{(\lambda)})-\frac{1}{2}R(g^{(\lambda)})g^{(\lambda)}_{\mu\nu}\Big)\frac{\partial g^{(\lambda)\mu\nu}}{\partial\lambda}\delta\lambda=0.$ (24)
Here, we consider an unbounded spacetime and neglect the boundary terms in the variation of the action. From the Einstein’s equation, we have
$\int d^{4}x\sqrt{-g^{(\lambda)}}\Big(R_{\mu\nu}(g^{(\lambda)})-\frac{1}{2}R(g^{(\lambda)})g^{(\lambda)}_{\mu\nu}\Big)\frac{\partial g^{(\lambda)\mu\nu}}{\partial\lambda}=2\lambda\int d^{4}x\sqrt{-g^{(\lambda)}}\hat{T}_{\mu\nu}(g^{(\lambda)})\frac{\partial g^{(\lambda)\mu\nu}}{\partial\lambda}=0.$ (25)
We now return to the action of Maxwell theory. For nonzero $\lambda$, taking the first order derivative with respect to $\lambda$ yields
$\frac{\text{d}}{\text{d}\lambda}S_{\text{Maxwell}}[g^{(\lambda)},A]$ $=\int d^{4}x\frac{\delta S_{\text{Maxwell}}}{\delta g^{\mu\nu}}\Big|_{g=g^{(\lambda)}}\frac{\partial g^{(\lambda)\mu\nu}}{\partial\lambda}$
$=\frac{1}{2}\int d^{4}x\sqrt{-g^{(\lambda)}}\hat{T}_{\mu\nu}(g^{(\lambda)})\frac{\partial g^{(\lambda)\mu\nu}}{\partial\lambda}=0.$ (26)
To demonstrate more explicitly that the Maxwell theory on the deformed metric is independent of the deformation parameter, we consider a static and spherically symmetric metric solution. Given the assumptions of time translation invariance and rotational invariance for the electromagnetic field tensor, it can possess at most two non-vanishing independent components [45, 12, 11]. The electromagnetic two-form can be formally written as
$F=A(r)\ dt\wedge dr+B(r)\sin\theta\ d\theta\wedge d\varphi.$ (27)
By employing the Maxwell equation, the coefficients of $F$ take the forms
$A(r)=-\frac{Q_{e}}{r^{2}},\ \ \ \ B(r)=Q_{m}.$ (28)
The metric solution is assumed to adopt the conventional static and spherically symmetric form,
$ds^{2}=-f^{(\lambda)}(r)dt^{2}+\frac{dr^{2}}{f^{(\lambda)}(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta\ d\varphi^{2}).$ (29)
The non-vanishing components of the inverse metric are $g^{(\lambda)tt}=-\frac{1}{f^{(\lambda)}(r)}$, $g^{(\lambda)rr}=f^{(\lambda)}(r)$, $g^{(\lambda)\theta\theta}=\frac{1}{r^{2}}$, and $g^{(\lambda)\varphi\varphi}=\frac{1}{r^{2}\sin^{2}\theta}$. It follows that
$F^{tr}=-A(r),\ \ \ F^{\theta\varphi}=\frac{B(r)}{r^{4}\sin\theta}.$ (30)
By plugging the components of the electromagnetic field and the metric into Einstein’s equation, we obtain the explicit form of $f^{(\lambda)}$, which is exactly the Reissner-Nordström solution [12],
$f^{(\lambda)}(r)=1-\frac{\lambda M}{4\pi r}+\frac{\lambda(Q_{e}^{2}+Q_{m}^{2})}{8\pi r^{2}}.$ (31)
Putting everything together, and using the definition of effective deformed field theory action, we find
$S_{\text{EDA}}[g^{(\lambda)},Q_{m},Q_{e}]=\int dtdrd\theta d\varphi\ \frac{\sin\theta(Q_{m}^{2}-Q_{e}^{2})}{r^{2}},$ (32)
which is independent of $\lambda$.
It is crucial to distinguish the above result carefully from that derived using a perturbative approach. Analogous to the scalar field case, we investigate the first-order non-local stress tensor deformation of Maxwell theory on a flat background,
$S^{(\lambda)}[\eta,A]$ $=S_{\text{Maxwell}}[\eta,A]+\frac{\lambda l^{-2}}{2}\int d^{4}x\sqrt{-\eta}\hat{T}_{\mu\nu}\frac{1}{-\hat{\Box}}(\hat{T}^{\mu\nu}-\frac{1}{2}\hat{T}^{\alpha}_{\alpha}\eta^{\mu\nu})+O(\lambda^{2})$
$=\int d^{4}x\sqrt{-\eta}\Big[-\frac{1}{4}F_{\alpha\beta}F^{\alpha\beta}+\frac{\lambda l^{-2}}{2}(F_{\mu\alpha}F_{\nu}{}^{\alpha}-\frac{1}{4}\eta_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta})\frac{1}{-\hat{\Box}}(F^{\mu}{}_{\alpha}F^{\nu\alpha}-\frac{1}{4}\eta^{\mu\nu}F_{\alpha\beta}F^{\alpha\beta})\Big]+O(\lambda^{2}).$ (33)
By introducing the metric redefinition
$g_{\mu\nu}$ $=\eta_{\mu\nu}+2\lambda l^{-2}\frac{1}{-\hat{\Box}}(F_{\mu\alpha}F_{\nu}{}^{\alpha}-\frac{1}{4}\eta_{\mu\nu}F_{\alpha\beta}F^{\alpha\beta})+O(\lambda^{2}),$
$g^{\mu\nu}$ $=\eta^{\mu\nu}-2\lambda l^{-2}\frac{1}{-\hat{\Box}}(F^{\mu}{}_{\alpha}F^{\nu\alpha}-\frac{1}{4}\eta^{\mu\nu}F_{\alpha\beta}F^{\alpha\beta})+O(\lambda^{2}),$ (34)
the deformed field theory action can be rewritten as
$S^{(\lambda)}$ $=-\frac{1}{4}\int d^{4}x\sqrt{-g}g^{\alpha\gamma}g^{\beta\delta}F_{\alpha\beta}F_{\gamma\delta}-\frac{l^{2}}{8\lambda}\int d^{4}x\sqrt{-\eta}(g^{\mu\nu}-\eta^{\mu\nu})\hat{\Box}(g_{\mu\nu}-\eta_{\mu\nu})+O(\lambda^{2}).$ (35)
This deformed action can be interpreted as a coupling between Maxwell theory on the deformed metric and linearized gravity. In contrast to the EFT approach, the linearized gravity sector here encompasses both the linearized Einstein–Hilbert term and the gauge-fixing term introduced in (Geometric realization of stress-tensor deformed field theory), (more precisely, the action should also incorporate additional ghost terms.) As a result, the first-order derivative of the deformed action with respect to $\lambda$ is non-vanishing.
@@
$S$ $=\hat{S}[\hat{\gamma}+\lambda h,\psi]+\frac{1}{2\lambda}S_{\text{grav}}[\hat{\gamma}+\lambda h]$
$=\hat{S}[\hat{\gamma},\psi]-\frac{\lambda}{2}\int d^{d}x\sqrt{\hat{\gamma}}h_{\mu\nu}\hat{T}^{\mu\nu}-\frac{1}{2}\sum_{k=1}^{\infty}\lambda^{k+1}\int d^{d}xh_{\mu\nu}\Big(\prod_{i=1}^{k}h_{\mu_{i}\nu_{i}}\Big)\frac{\partial^{k}}{\prod_{i=1}^{k}\partial\hat{\gamma}_{\mu_{i}\nu_{i}}}\Big(\sqrt{\hat{\gamma}}\hat{T}^{\mu\nu}\Big)$
@@
$\mathcal{F}^{(k)}$ $=\mathcal{F}^{(k)}(\hat{\gamma}_{\mu\nu},h_{\mu\nu},\hat{\nabla}_{\rho}h_{\mu\nu},\hat{\nabla}_{\rho}\hat{\nabla}_{\sigma}h_{\mu\nu})\sim\lambda^{k}h^{k},$
@@
The EOM for the auxiliary field is
$-\lambda l^{-2}\Big[{\hat{T}^{\mu\nu}+\sum_{k=1}^{\infty}\Big(\lambda^{k}\prod_{i=1}^{k}h_{\mu_{i}\nu_{i}}\Big)\frac{1}{\sqrt{\hat{\gamma}}}\Big(\frac{\partial^{k}}{\prod_{i=1}^{k}\partial\hat{\gamma}_{\mu_{i}\nu_{i}}}(\sqrt{\hat{\gamma}}\hat{T}^{\mu\nu})+k\frac{\partial^{k}}{\partial\hat{\gamma}_{\mu\nu}\prod_{i=2}^{k}\partial\hat{\gamma}_{\mu_{i}\nu_{i}}}(\sqrt{\hat{\gamma}}\hat{T}^{\mu_{1}\nu_{1}})\Big)}\Big]$
$\quad+\sum_{k=0}^{\infty}\tilde{\mathcal{F}}^{(k)\mu\nu}=0,$ (38)
@@
$\tilde{\mathcal{F}}^{(k)\mu\nu}(x)=\frac{1}{\sqrt{\hat{\gamma}(x)}}\frac{\delta}{\lambda\delta h_{\mu\nu}(x)}\int d^{d}x^{\prime}\sqrt{\hat{\gamma}}\mathcal{F}^{(k+1)}\sim\lambda^{k}h^{k}.$ (39)
@@
$\tilde{\mathcal{F}}^{(0)\mu\nu}$ $=\frac{1}{2}\hat{R}\hat{\gamma}^{\mu\nu}-\hat{R}^{\mu\nu},$
$\tilde{\mathcal{F}}^{(1)\mu\nu}$ $=\frac{\lambda}{2}\hat{\Box}h^{\mu\nu}-\frac{\lambda}{4}\hat{\gamma}^{\mu\nu}\hat{\Box}h^{\alpha}_{\alpha}+\frac{\lambda}{2}\Big[\hat{R}^{\mu\rho\nu\sigma}+\hat{R}^{\nu\rho\mu\sigma}+\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}+\hat{R}^{\nu\rho}\hat{\gamma}^{\mu\sigma}$
$\quad-\hat{R}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-\hat{R}^{\rho\sigma}\hat{\gamma}^{\mu\nu}-\frac{1}{2}\hat{R}(2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\Big]h_{\rho\sigma}.$ (40)
@@
$h^{*}_{[k]\mu\nu}=0,\text{ for }k<0.$ (42)
Such a saddle point corresponds to the seed theory with a non-local stress tensor deformation, and the form of the deformation can be solved order by order using the perturbation method. Below, we proceed to calculate the deformation up to the second order in $\lambda$. One can easily find that
$\tilde{\mathcal{F}}^{(2)\mu\nu}$ $=\frac{\lambda^{2}}{8}\Big(4\hat{\nabla}_{\rho}h^{\mu\nu}\hat{\nabla}_{\sigma}h^{\rho\sigma}-8\hat{\nabla}_{\rho}h^{\mu\rho}\hat{\nabla}_{\sigma}h^{\nu\sigma}-2h^{\mu\nu}\hat{\nabla}_{\sigma}\hat{\nabla}_{\rho}h^{\rho\sigma}+8h^{\nu\rho}\hat{\nabla}_{\sigma}\hat{\nabla}_{\rho}h^{\mu\sigma}+4h^{\rho\sigma}\hat{\nabla}_{\sigma}\hat{\nabla}_{\rho}h^{\mu\nu}+2h^{\mu\nu}\hat{\Box}h^{\rho}_{\rho}$
@@
$h^{*}_{\mu\nu}(x)$ $=-2t_{\mu\nu}(x)+4\lambda l^{-2}\int d^{d}y\sqrt{\hat{\gamma}(y)}G_{\mu\nu\eta\xi}(x,y)\Big(\delta^{\eta}_{\alpha}\delta^{\xi}_{\beta}-\frac{1}{d-2}\hat{\gamma}^{\eta\xi}(y)\hat{\gamma}_{\alpha\beta}(y)\Big)\Big[M^{\alpha\beta\rho\sigma}t_{\rho\sigma}$
@@
The perturbation method can be employed to compute stress tensor deformations associated with more general higher-order gravitational actions. An important example is the Einstein gravity coupled with the Gauss-Bonnet term,
$S_{G}$ $=S_{\text{EH}}+\alpha S_{\text{GB}}$
$=l^{2}\int d^{d}x\sqrt{g}\Big[R+\alpha l^{2}(R^{2}-4R^{\mu\nu}R_{\mu\nu}+R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma})\Big].$ (47)
For a saddle point satisfying condition $h^{*}=O(\lambda^{0})$ to exist, the background metric must satisfy the Einstein equation
$\hat{G}_{\mu\nu}+2\alpha l^{2}(\hat{R}\hat{R}_{\mu\nu}-2\hat{R}_{\mu\alpha}\hat{R}^{\alpha}{}_{\nu}-2\hat{R}^{\alpha\beta}\hat{R}_{\mu\alpha\nu\beta}+\hat{R}_{\mu}{}^{\alpha\beta\gamma}\hat{R}_{\nu\alpha\beta\gamma}-\frac{1}{4}\hat{\mathcal{L}}_{\text{GB}}\hat{\gamma}_{\mu\nu})=0.$ (48)
The simplest choice is the flat background metric $\gamma_{\mu\nu}=\eta_{\mu\nu}$. Expanding the Gauss-Bonnet term in powers of $h_{\mu\nu}$, we have
$S_{\text{GB}}=l^{4-D}\sum_{k=0}^{\infty}\int d^{D}x\sqrt{\eta}\,\mathcal{R}^{(k)}_{\text{GB}}$, (49)
with
$\mathcal{R}^{(0)}_{\text{GB}}$ $=\mathcal{R}^{(1)}_{\text{GB}}=0,$
$\mathcal{R}^{(2)}_{\text{GB}}$ $=\frac{\lambda^{2}}{2}(8\partial_{\mu}\partial^{\rho}h^{\mu\nu}\partial_{\rho}\partial_{\nu}h-4\partial_{\nu}\partial_{\mu}h\partial^{2}h^{\mu\nu}-4\partial_{\nu}\partial_{\mu}h^{\mu\nu}\partial^{2}h-4\partial_{\mu}\partial^{\rho}h^{\mu\nu}\partial_{\sigma}\partial_{\nu}h_{\rho}^{\sigma}$
$\quad+2\partial_{\nu}\partial_{\mu}h^{\mu\nu}\partial_{\sigma}\partial_{\rho}h^{\rho\sigma}+2\partial_{\nu}\partial_{\mu}h_{\rho\sigma}\partial^{\sigma}\partial^{\rho}h^{\mu\nu}-4\partial_{\mu}\partial^{\rho}h^{\mu\nu}\partial_{\sigma}\partial_{\rho}h_{\nu}^{\sigma}+8\partial_{\mu}\partial^{\rho}h^{\mu\nu}\partial^{2}h_{\nu\rho}$
$\quad-2\partial_{\sigma}\partial_{\nu}h_{\mu\rho}\partial^{\sigma}\partial^{\rho}h^{\mu\nu}-2\partial_{\rho}\partial_{\nu}h\partial^{\rho}\partial^{\nu}h+2\partial^{2}h\partial^{2}h-2\partial_{\nu}\partial_{\rho}h_{\mu\sigma}\partial^{\sigma}\partial^{\rho}h^{\mu\nu}$
$\quad-2\partial^{2}h^{\mu\nu}\partial^{2}h_{\mu\nu}+2\partial_{\sigma}\partial_{\rho}h_{\mu\nu}\partial^{\sigma}\partial^{\rho}h^{\mu\nu}).$ (50)
When the background metric is flat, $\mathcal{R}^{(2)}_{\text{GB}}$ is a total derivative term. Therefore, the inclusion of the Gauss-Bonnet term does not change the first-order form of the stress tensor deformation (12).